OpenStudy (anonymous):

Find the product of (2x − 5y)^2. A) 4x^2 + 20xy + 25y^2 B) 4x^2 − 20xy + 25y^2 C) 4x^2 − 25y^2 D) 4x^2 + 25y^2

OpenStudy (anonymous):

@Notamathgenius

OpenStudy (anonymous):

@dmezzullo

OpenStudy (dmezzullo):

what is 4x4? what is 5x5?

OpenStudy (anonymous):

4x^2-20xy+25y^2

OpenStudy (anonymous):

16 and 25

OpenStudy (lopus):

is simple is B only is (2x-5y)*(2x-5y)

OpenStudy (anonymous):

thats the answer

OpenStudy (dmezzullo):

Wait wrong thing, lemme redo it

OpenStudy (anonymous):

the answe is D

OpenStudy (dmezzullo):

ok so 2x2 is ?? and 5x5 is ??

OpenStudy (dmezzullo):

@bahupepe Do not jus give the answer out explain it also this is a site to learn. Welcome to OPenstudy!! also

OpenStudy (anonymous):

So its D?

OpenStudy (lopus):

OMG

OpenStudy (dmezzullo):

Yes it is D, But in the future guys plz explain the problem.

OpenStudy (anonymous):

Find the product of (x + 3)^2. x^2 + 6x + 9 x^2 − 6x + 9 x^2 − 9 x^2 + 9

OpenStudy (anonymous):

okay (2x-5y)(2x-5y) 4x^2-10xy-10xy+25y^2 4x^2-20xy+25y^2

OpenStudy (dmezzullo):

Ok so x times x ?? 3x3 ??

OpenStudy (anonymous):

3x3 = 9 x times x = idk

OpenStudy (dmezzullo):

ok so x times x would jus be x^2 bc u are jus timesing it by itself.

OpenStudy (dmezzullo):

So then wut do u think the answer is?

OpenStudy (lopus):

(x+3)*(x+3)= (x^2+3x+3x+9) you're joke

OpenStudy (anonymous):

oh okay. so am i adding or subtracting?

OpenStudy (dmezzullo):

Adding.

OpenStudy (lopus):

x*X=x^2

OpenStudy (anonymous):

So D?

OpenStudy (dmezzullo):

Yes.